# Published in the proceedings of the 39th Conference on Neural Information Processing Systems (NeurIPS 2025) Workshop: Scaling Environments for Agents (SEA). Additionally accepted for presentation in NeurIPS 2025 Workshop: Embodied World Models for Decision Making

### v2

@@ 2 Related work
The evolution from flat, 2D cultures to 3D neural organoids marks a profound leap in biological fidelity. Organoids more closely recapitulate the complex cytoarchitecture, cellular diversity, and developmental trajectories of the human brain, offering a far richer substrate for investigating the network-level phenomena that underpin complex decision-making and the formation of internal models (Lancaster et al., 2013). Recent work has successfully demonstrated that these organoid-based systems can learn to perform tasks, such as playing a simplified version of the game Pong, by harnessing their inherent plasticity (Kagan et al., 2022).
@@ 3.1 Multi-electrode array interface
All proposed experiments are built upon a multi-electrode array (MEA) platform, which provides the critical bidirectional interface for embodying the biological agent. This platform must be capable of concurrent, high-fidelity stimulation and recording from the neural culture. For simplicity and reproducibility, we assume a setup with at least four distinct electrode groups (A, B, C, D), which can be spatially organized across the array. Each group can consist of one or more electrodes. Groups A and B are designated for recording spontaneous and evoked neural activity, serving as the agent’s motor output channels. Groups C and D are dedicated to delivering patterned electrical stimuli, functioning as the agent’s sensory input channels. This architecture forms the physical link between the virtual environment and the biological substrate.
@@ 3.2 Biological substrates and embodiment
While this paper focuses on human neural organoids due to their advanced structural and cellular complexity that emulates aspects of the developing human brain, the experimental framework we propose is fundamentally substrate-agnostic. The core principles of sensory encoding, action decoding, and feedback-driven learning can be readily adapted to a variety of biological preparations. These include simpler neural spheroids, traditional 2D neuronal cultures, more structured brain-on-a-chip platforms, or even non-neural tissues that exhibit adaptive plasticity. The essential requirement is a biological system that can undergo experience-dependent changes and can be integrated into a closed-loop control system via an MEA for real-time stimulation and recording. This inherent flexibility allows for powerful comparative studies across different scales of biological organization, from dissociated networks to complex, multi-layered tissues.
@@ 3.3 Inducing world model formation via predictive coding and reinforcement
Reward (Model Confirmation). A reward signal, delivered upon achieving a desirable state (e.g., capturing prey), is implemented as a predictable, low-entropy electrical stimulus. A concrete implementation involves delivering a consistent, low-frequency sinusoidal wave across all stimulation electrodes. This spatially and temporally predictable sensory input minimizes surprise, serving as a confirmation signal that reinforces the agent’s preceding policy and strengthens its internal model of successful state-action sequences. An alternative biological implementation involves the targeted delivery of dopamine via UV light-controlled uncaging, directly activating the brain’s reward pathways known to be crucial for inducing long-term potentiation (Schultz et al., 1997).
Punishment (Model Violation). Conversely, a punishment signal, corresponding to an undesirable outcome (e.g., entering an aversive zone), is delivered as an unpredictable, high-entropy stimulus. This is implemented by applying a white-noise electrical signal, characterized by random amplitude and frequency, to a randomly selected subset of stimulation electrodes. This chaotic and surprising sensory input functions as a powerful error signal, indicating a violation of the agent’s predictive model. The objective is to drive the network to update its policy to avoid states and actions that lead to such high-surprise, model-violating outcomes.
@@ 3.4 Evaluating the embodied world model: From task performance to synaptic correlates
At the functional network level, we can probe synaptic efficacy by measuring field excitatory postsynaptic potentials (fEPSPs) before and after training epochs. By delivering a controlled test pulse to a stimulation electrode and recording the evoked potential in a downstream population, the slope of the fEPSP serves as a robust proxy for synaptic strength. A sustained increase in this slope post-training is the canonical electrophysiological signature of long-term potentiation (LTP), while a decrease signifies long-term depression (LTD). This provides a direct, quantifiable link between the agent’s interactive experience and the Hebbian plasticity mechanisms thought to underlie learning and memory (Bliss and Lomo, 1973; Malenka and Bear, 2004).
To resolve how neural populations encode task variables and represent the world model at a finer scale, we can employ optical imaging. In organoids engineered to express Genetically Encoded Calcium Indicators (GECIs) like GCaMP (Chen et al., 2013), two-photon microscopy enables longitudinal tracking of the activity of hundreds of individual neurons. This technique allows us to observe the evolution of network dynamics as learning progresses, potentially identifying the emergence of stable neural ensembles that represent specific states, actions, or environmental features. Changes in the synchrony, firing rate, and spatial organization of these calcium transients can elucidate how the agent’s internal model is formed, stored, and refined through goal-directed interaction.
Finally, at the conclusion of an experiment, post-hoc molecular analyses can reveal the structural correlates of the observed functional changes. Immunohistochemistry can be used to visualize the molecular machinery of plasticity. For example, staining for the differential trafficking of AMPA and NMDA receptor subunits can expose changes in synaptic receptor density, a key mechanism underlying the expression of LTP and LTD (Malenka and Bear, 2004). Furthermore, using antibodies against plasticity-related phosphorylated proteins, such as pCaMKII or pGluA1, provides a molecular snapshot of synapses that were recently potentiated, thereby linking task-relevant neural activity to specific structural modifications (Heo et al., 2023). By integrating these electrophysiological, optical, and molecular benchmarks, we can construct a comprehensive portrait of learning that bridges behavioral performance with the fundamental principles of neural adaptation, offering a ground-truth look into the learned world model.
@@ 4.1.1 Protocol description
4.1.2 Environmental scaling and long-horizon planning
  • Dimensionality: The task’s complexity can be significantly increased by scaling the world from a 1D line to a 2D grid or a 3D volume. A 2D environment requires a more sophisticated world model representing an (x, y) state, which could be encoded using separate electrode groups or frequency channels for each axis. The action space must expand to four cardinal directions, demanding a more complex motor decoding scheme, such as using four distinct recording sites or interpreting more complex spatio-temporal firing patterns. A pseudocode implementation is provided in Listing 5.
  • Zone Complexity and Planning: The simple binary safe/aversive partition can be replaced with more complex spatial arrangements that require rudimentary planning. For example, the agent could be rewarded for navigating a narrow, winding path or for solving a simple maze defined by aversive boundaries. This challenges the agent to build a world model that represents not just a single boundary, but an extended spatial policy. See Listing 6.
  • Dynamic Environments and Adaptation: To probe the adaptability of the learned world model, the environment can be made non-stationary. The boundary of the safe zone could slowly shift over the course of an experiment, or the goal location in a maze could change after a block of successful trials. This forces the agent to extinguish a previously learned policy and acquire a new one, providing a direct assay for the flexibility and update speed of its internal model (see Listing 7).
4.2 Environment 2: Goal-seeking in a predator-prey scenario
4.2.1 Protocol description
  1. 1.
  2. 2.
  3. 3.
4.2.2 Environmental scaling and multi-agent systems
  • Dimensionality and Search Strategy: Scaling the world to 2D or 3D transforms the task from a simple directional problem into a genuine search problem. The agent must learn an efficient search strategy (e.g., patterned, memory-guided, or random) to locate the prey in a vastly larger state space. This places significant demands on the world model’s capacity to integrate sensory information over extended time horizons to infer the target’s location. See Listing 8.
  • Introduction of an Adversary: A second mobile entity, an adversary, can be introduced, creating a multi-objective task: approach prey while avoiding the adversary. The adversary’s location would be encoded on a separate sensory channel (e.g., via a distinct stimulation frequency). Its behavior could range from a random walk to actively tracking the agent, requiring the agent’s world model to account for multiple dynamic entities with different behavioral characteristics. See Listing 9.
  • Multi-Organoid Systems: The framework can be extended to create a simple competitive ecosystem. One organoid could control the prey and a second could control the predator, allowing their policies to co-evolve. This can be further extended to a linear food chain (Organoid A is prey to B, B is prey to C) or a cyclical model, providing a novel platform to study the emergence of complex, hierarchical strategies in a multi-agent biological system. A pseudocode implementation for a two-organoid system is available in Listing 10.
4.3 Environment 3: Dynamic interception task (Pong)
4.3.1 Protocol description
  1. 1.
  2. 2.
  3. 3.
  4. 4.
4.3.2 Environmental scaling and generalization
  • Generalization to Other Games: The core paradigm of Pong can be adapted to other classic arcade games to probe different cognitive functions and world modeling capabilities. For instance, Breakout would require the agent’s world model to incorporate not just the ball and paddle, but also a static array of targets, demanding precise spatial targeting to aim the ball. Space Invaders would introduce the complexities of dodging projectiles while timing offensive actions, requiring the world model to track multiple independent agents. A pseudocode implementation for Breakout is provided in Listing 11.
  • Competitive Multi-Organoid Systems: Pong provides an ideal testbed for multi-agent competition. Two distinct organoid-MEA systems can be linked to the same game instance, each controlling one of the paddles. This setup facilitates the direct study of competitive co-adaptation, where each biological agent must build a world model that includes the policy of its opponent. The evolution of rally lengths, shot patterns, and win/loss rates would serve as powerful metrics for inter-agent learning and strategic model formation (see Listing 12).
Figure 3: Diagram of the Pong environment. The organoid receives complex, multi-modal sensory input about the ball’s position (rate-coded distance, spatially-coded location) and must learn a predictive world model of the ball’s trajectory to control its paddle and achieve a successful interception. The red arrow shows an example trajectory for the orange ball: starting at the top, bouncing off of walls and the agent’s paddle, bouncing off additional walls, and finally passing the agent.
To overcome the limitations of manual experimental design and to systematically explore the vast parameter space of environments and training protocols, we propose an automated framework where an LLM functions as a meta-controller or AI Dungeon Master. This approach directly aligns with the theme of leveraging generative models and LLMs, applying the LLM not as the embodied agent, but as a tool to orchestrate the generation of the agent’s training world and curriculum. This creates a high-throughput system for discovering optimal training strategies and environmental configurations that might be non-obvious to human researchers. A pseudocode implementation of this meta-learning loop is provided in Listing 4.
4.4.1 Framework description
  1. 1.
  2. 2.
    Protocol Generation: The LLM processes the prompt and generates a novel experimental protocol. This can be done in two primary modes: generating a structured JSON object with specific experimental variables to be plugged into a template script, or generating an entire Python script from scratch that defines novel experimental logic and curriculum stages. This generative capacity allows the LLM to move beyond simple parameter tuning to design entirely new tasks. Detailed examples are provided in Appendix A.3.
  3. 3.
  4. 4.
  5. 5.
5 Discussion
5.1 Broader implications and future work
5.2 Limitations and challenges
Despite its significant promise, this approach is not without its challenges. The biological substrate introduces inherent variability, limited lifespans, and developmental stochasticity, making reproducibility a central concern that must be addressed with robust experimental design and statistical analysis. The organoid itself remains a profound black box; while we can measure network-level plasticity, deciphering the specific computational algorithms it implements to learn and represent its world model is a formidable open problem. Furthermore, the information bandwidth of current MEA technology, while improving, is still orders of magnitude lower than that of biological nervous systems, potentially constraining the complexity of the world models that can be learned.
6 Conclusion
Acknowledgments and Disclosure of Funding
We thank FinalSpark for proposing the idea of LLM based automated research.
Appendix A Technical appendices and supplementary material
A.1 Additional pseudocode for initial environments and automation
This section provides the base pseudocode for the Pong environment and the LLM-driven automation framework described in the main text.
@@ A.1 Additional pseudocode for initial environments and automation
Listing 2: Pseudocode for predator-prey task
Initialize game state and MEA interface
game = PongGame()
Main game loop
while game.is_active:
# 1. Update game state
game.update_ball_position()
ball_x, ball_y = game.get_ball_position()
# 2. State representation
encode_y_pos_rate_coded(groups_C_D, ball_y)
encode_x_pos_spatial(groups_C_D, ball_x)
# 3. Action decoding
spikes_A = record_spikes(group_A, window=10_ms)
spikes_B = record_spikes(group_B, window=10_ms)
if spikes_A > spikes_B:
game.move_paddle_up()
elif spikes_B > spikes_A:
game.move_paddle_down()
# 4. Feedback
if game.check_collision():
deliver_reward_predictable(groups_C_D)
game.ball_bounce()
elif game.check_miss():
deliver_punishment_unpredictable(groups_C_D)
game.reset_ball()
Listing 3: Pseudocode for pong task
Initialize dataset and load base prompt template
dataset = initialize_database()
prompt_template = load_base_prompt("config/llm_prompt.txt")
Main automation loop
for iteration in range(NUM_ITERATIONS):
# 1. Formulate prompt with historical data
current_organoid_state = get_organoid_baseline()
# 2. LLM generates a new experimental script
generated_script = call_llm_api(prompt)
# 3. Validate the generated script for safety and syntax
# 4. Execute or log error
if is_valid:
# Run the experiment on the MEA platform
[…]
# Store the script and its outcomes in the dataset
else:
# Log the invalid script and the reason for failure
# 5. Periodically refine the LLM’s instructions
if (iteration + 1) % BATCH_SIZE == 0:
prompt_template = refine_prompt_based_on_results(dataset)
# Optional: Trigger fine-tuning if sufficient data is collected
# if len(dataset) > FINETUNE_THRESHOLD:
# launch_finetuning_job(dataset)
Listing 4: Pseudocode for LLM-driven experiment automation
A.2 Pseudocode for scaled environments
This section provides pseudocode implementations for the scaling variations. The algorithms are grouped by the base environment from which they are extended, providing concrete examples of how task complexity can be systematically increased.
A.2.1 Scaling the conditional avoidance environment
The algorithms below correspond to the scaling variations for the Conditional Avoidance task, including extensions to 2D space, maze navigation, and dynamic (non-stationary) conditions.
Initialize a 10x10 grid environment
GRID_SIZE = 10
Assume 4 recording sites (A_up, A_down, B_left, B_right)
initialize_platform(groups_A_up, A_down, B_left, B_right, C, D)
while experiment_running:
# 1. State representation (2D)
# 2. Action decoding (4-way movement)
spikes_up = record_spikes(group_A_up, window=T_ms)
spikes_down = record_spikes(group_A_down, window=T_ms)
spikes_left = record_spikes(group_B_left, window=T_ms)
spikes_right = record_spikes(group_B_right, window=T_ms)
# Update position based on decoded action
x, y = agent_pos
if action == ’UP’: y = min(GRID_SIZE - 1, y + 1)
if action == ’DOWN’: y = max(0, y - 1)
if action == ’LEFT’: x = max(0, x - 1)
if action == ’RIGHT’: x = min(GRID_SIZE - 1, x + 1)
agent_pos = (x, y)
# 3. Feedback based on 2D zone
if agent_pos[0] in aversive_zone_x:
punishment_amplitude = calculate_amplitude(agent_pos[0])
else:
Listing 5: Pseudocode for 2D conditional avoidance
Define a maze as a grid where 1=path, 0=wall
maze_layout = [
[1, 1, 1, 0, 1],
[0, 1, 0, 0, 1],
[1, 1, 1, 1, 1],
[1, 0, 1, 0, 1],
[1, 1, 1, 0, 1]
start_pos, goal_pos = (0, 0), (4, 4)
agent_pos = start_pos
while experiment_running:
# State and Action decoding remain similar to 2D avoidance
action = decode_4way_movement(groups_A_B)
next_pos = calculate_next_pos(agent_pos, action)
# 3. Feedback and state transition based on maze structure
x, y = next_pos
if maze_layout[y][x] == 1: # Valid path
agent_pos = next_pos
if agent_pos == goal_pos:
deliver_reward(groups_C_D)
agent_pos = start_pos # Reset to start
else: # Hit a wall
deliver_punishment(groups_C_D)
# Agent does not move if it tries to enter a wall
Listing 6: Pseudocode for maze navigation task
Initialize environment with a moving boundary
agent_position = 4
safe_zone_boundary = 5 # Initially, positions 1-5 are safe
boundary_shift_interval = 100 # Timesteps before boundary moves
timestep = 0
while experiment_running:
# 1. State representation & 2. Action decoding (same as 1D)
encode_position_as_stimulus(groups_C_D, agent_position)
# 3. Feedback based on the DYNAMIC boundary
if agent_position <= safe_zone_boundary:
deliver_reward(groups_C_D)
else:
deliver_punishment(groups_C_D)
# 4. Update the environment itself
timestep += 1
if timestep % boundary_shift_interval == 0:
if random.choice([True, False]):
safe_zone_boundary = max(1, safe_zone_boundary - 1)
else:
safe_zone_boundary = min(safe_zone_boundary + 1, 8)
Listing 7: Pseudocode for dynamic avoidance task
A.2.2 Scaling the predator-prey environment
These examples illustrate extensions to the predator-prey scenario. This includes increasing dimensionality, introducing an adversary, and creating a multi-agent competitive system.

### v3

@@ 2 Related work
The evolution from flat, 2D cultures to 3D neural organoids marks a profound leap in biological fidelity (Lancaster et al., 2013). Organoids leverage the principles of developmental self-organization (Zhang et al., 2022) to recapitulate the complex cytoarchitecture, cellular diversity, and developmental trajectories of the human brain. This methodology has produced both unguided cerebral organoids, which stochastically generate multiple brain regions to model inter-regional interactions (Lancaster and Knoblich, 2014), and guided organoids, which use morphogens to direct differentiation toward specific brain regions (e.g., cortex) for improved reproducibility (Shaker et al., 2022; Pavlinek et al., 2025). This offers a far richer substrate for investigating the network-level phenomena that underpin complex decision-making and the formation of internal models. Recent work has successfully demonstrated that these organoid-based systems can learn to perform tasks, such as playing a simplified version of the game Pong, by harnessing their inherent plasticity (Kagan et al., 2022).
@@ 3.1 Multi-electrode array interface
All proposed experiments are built upon a multi-electrode array (MEA) platform, which provides the critical bidirectional interface for embodying the biological agent. This platform must be capable of concurrent, high-fidelity stimulation and recording from the neural culture. For simplicity and reproducibility, we assume a setup with at least four distinct electrode groups (A, B, C, D), which can be spatially organized across the array. Each group can consist of one or more electrodes. Groups A and B are designated for recording spontaneous and evoked neural activity, serving as the agent’s motor output channels. Groups C and D are dedicated to delivering patterned electrical stimuli, functioning as the agent’s sensory input channels. However, the scaling of environments and the fine-grained evaluation of the world model ideally require more advanced interfaces. These include high-density CMOS-based MEAs (HD-MEAs), which offer thousands of electrodes for high-resolution field potential imaging (Yokoi et al., 2025), or flexible mesh and 3D-pillar MEAs that achieve superior volumetric integration and signal quality with 3D tissues (McDonald et al., 2023). This architecture forms the physical link between the virtual environment and the biological substrate.
@@ 3.2 Biological substrates and embodiment
The primary biological substrate for this work is the human neural organoid, chosen for its structural and cellular complexity (Lancaster et al., 2013). The success of any training protocol is critically contingent on the developmental maturity and long-term viability of these organoids.
Maturation and Culture Requirements: Neural organoids require a protracted in vitro culture period, typically a minimum of 60-90 days, to develop the cellular diversity (including neurons, astrocytes, and other glia (Porciúncula et al., 2021)) and functional network activity, such as synchronized bursting, required for learning (Fair et al., 2020; Giandomenico et al., 2021). During this extended maturation, organoids must be maintained in a dynamic culture system, such as on an orbital shaker or in a spinning bioreactor. This is essential to enhance nutrient and oxygen diffusion, preventing the formation of a hypoxic or necrotic core that compromises tissue health and experimental reproducibility (Zhang et al., 2022). Long-term viability is further supported by specialized, serum-free media formulations (e.g., BrainPhys-based media) supplemented with neurotrophic factors like BDNF and GDNF (Zhang et al., 2022; Fair et al., 2020).
Protocol Choice and Reproducibility: The experiments proposed can utilize organoids from either guided or unguided differentiation protocols. Guided protocols, which use exogenous morphogens to direct differentiation towards a specific regional identity like the dorsal forebrain (Shaker et al., 2022), are preferred for tasks requiring high reproducibility and lower inter-organoid variability. Unguided protocols, which rely on intrinsic self-organization, generate a more complex and heterogeneous mix of brain regions (Lancaster and Knoblich, 2014), but this stochasticity presents a significant challenge for controlled learning experiments. Therefore, a semi-guided protocol (Fitzgerald et al., 2024) or the use of guided cortical organoids represents the most promising balance of complexity and consistency.
Integration with the MEA: For all experiments, healthy, mature organoids must be selected and carefully plated onto the MEA surface, which is pre-treated with adhesion-promoting molecules like Poly-D-Lysine and Laminin (Hales et al., 2010). A stable, high-quality electrical interface is achieved by using surface tension to gently attach the organoid (Yokoi et al., 2025) and allowing the system to acclimatize at 37°C and 5% CO2 for at least 45-90 minutes before recording to establish a stable baseline (Fair et al., 2020). While this paper focuses on whole organoids, an alternative for higher-throughput screening of the LLM-generated protocols would be to use 2.5D dissociated organoid cultures, which sacrifice 3D architecture for improved homogeneity (Pavlinek et al., 2025).
@@ 3.3 Inducing world model formation via predictive coding and reinforcement
Reward (Model Confirmation). A reward signal, delivered upon achieving a desirable state (e.g., capturing prey), is implemented as a predictable, low-entropy electrical stimulus. A concrete implementation involves delivering a consistent, low-frequency sinusoidal wave across all stimulation electrodes. This spatially and temporally predictable sensory input minimizes surprise, serving as a confirmation signal that reinforces the agent’s preceding policy and strengthens its internal model of successful state-action sequences. An alternative biological implementation involves the targeted delivery of dopamine via UV light-controlled uncaging. This is achieved by perfusing the culture with a biologically inactive ’caged’ dopamine compound; a focused pulse of light then cleaves the caging group, releasing active dopamine at a precise spatiotemporal location (Ellis-Davies, 2019). This directly activates the brain’s reward pathways, which are known to encode reward prediction errors (Schultz, 1998) and are crucial for inducing long-term potentiation (Schultz et al., 1997).
Punishment (Model Violation). Conversely, a punishment signal, corresponding to an undesirable outcome (e.g., entering an aversive zone), is delivered as an unpredictable, high-entropy stimulus. This is implemented by applying a white-noise electrical signal, characterized by random amplitude and frequency, to a randomly selected subset of stimulation electrodes. This chaotic and surprising sensory input functions as a powerful error signal, indicating a violation of the agent’s predictive model. The objective is to drive the network to update its policy to avoid states and actions that lead to such high-surprise, model-violating outcomes, providing a direct test of computational theories like the Free Energy Principle (Mazzaglia et al., 2022).
@@ 3.4 Evaluating the embodied world model: From task performance to synaptic correlates
At the functional network level, we can probe synaptic efficacy by measuring field excitatory postsynaptic potentials (fEPSPs) before and after training epochs. By delivering a controlled test pulse to a stimulation electrode and recording the evoked potential in a downstream population, the slope of the fEPSP serves as a robust proxy for synaptic strength. A sustained increase in this slope post-training is the canonical electrophysiological signature of long-term potentiation (LTP), while a decrease signifies long-term depression (LTD). This plasticity can be actively probed by applying specific induction protocols, such as high-frequency tetanic stimulation (e.g., 100 Hz) to induce LTP or prolonged low-frequency stimulation (e.g., 1 Hz) to induce LTD (Nabavi et al., 2014). This provides a direct, quantifiable link between the agent’s interactive experience and the Hebbian plasticity mechanisms thought to underlie learning and memory (Bliss and Lomo, 1973; Malenka and Bear, 2004).
To resolve how neural populations encode task variables and represent the world model at a finer scale, we can employ optical imaging. In organoids engineered to express Genetically Encoded Calcium Indicators (GECIs) like GCaMP (Chen et al., 2013), often delivered via AAV transduction (Giandomenico et al., 2021), two-photon microscopy enables longitudinal tracking of the activity of hundreds of individual neurons. This approach is ideally performed simultaneously with MEA recording to correlate cellular-level calcium transients with network-level electrical events (Irwin and Allen, 2013). This technique allows us to observe the evolution of network dynamics as learning progresses, potentially identifying the emergence of stable neural ensembles that represent specific states, actions, or environmental features. Changes in the synchrony, firing rate, and spatial organization of these calcium transients can elucidate how the agent’s internal model is formed, stored, and refined through goal-directed interaction.
Finally, at the conclusion of an experiment, post-hoc molecular analyses can reveal the structural correlates of the observed functional changes. Immunohistochemistry can be used to visualize the molecular machinery of plasticity. For example, staining for the differential trafficking of AMPA and NMDA receptor subunits can expose changes in synaptic receptor density, a key mechanism underlying the expression of LTP and LTD (Malenka and Bear, 2004). This can be complemented by staining for presynaptic (e.g., Synapsin-1) and postsynaptic (e.g., PSD-95) markers to quantify changes in synapse density (Yakoub and Sadek, 2019), general cell-type markers (e.g., MAP2 for neurons, GFAP for astrocytes) to assess culture health, and activity-dependent immediate early genes like c-Fos to identify the specific neuronal ensembles that were recently active (Heo et al., 2023). By integrating these electrophysiological, optical, and molecular benchmarks, we can construct a comprehensive portrait of learning that bridges behavioral performance with the fundamental principles of neural adaptation, offering a ground-truth look into the learned world model.
@@ 4.1.1 Protocol description
4.1.2 Environmental scaling and long-horizon planning
  • Dimensionality: The task’s complexity can be significantly increased by scaling the world from a 1D line to a 2D grid or a 3D volume. A 2D environment requires a more sophisticated world model representing an (x, y) state, which could be encoded using separate electrode groups or frequency channels for each axis. The action space must expand to four cardinal directions, demanding a more complex motor decoding scheme, such as using four distinct recording sites or interpreting more complex spatio-temporal firing patterns. A pseudocode implementation is provided in Listing 2.
    Initialize a 10x10 grid environment
    GRID_SIZE = 10
    Assume 4 recording sites (A_up, A_down, B_left, B_right)
    initialize_platform(groups_A_up, A_down, B_left, B_right, C, D)
    while experiment_running:
    # 1. State representation (2D)
    # 2. Action decoding (4-way movement)
    spikes_up = record_spikes(group_A_up, window=T_ms)
    spikes_down = record_spikes(group_A_down, window=T_ms)
    spikes_left = record_spikes(group_B_left, window=T_ms)
    spikes_right = record_spikes(group_B_right, window=T_ms)
    # Update position based on decoded action
    x, y = agent_pos
    if action == ’UP’: y = min(GRID_SIZE - 1, y + 1)
    if action == ’DOWN’: y = max(0, y - 1)
    if action == ’LEFT’: x = max(0, x - 1)
    if action == ’RIGHT’: x = min(GRID_SIZE - 1, x + 1)
    agent_pos = (x, y)
    # 3. Feedback based on 2D zone
    if agent_pos[0] in aversive_zone_x:
    punishment_amplitude = calculate_amplitude(agent_pos[0])
    else:
    Listing 2: Pseudocode for 2D conditional avoidance
  • Zone Complexity and Planning: The simple binary safe/aversive partition can be replaced with more complex spatial arrangements that require rudimentary planning. For example, the agent could be rewarded for navigating a narrow, winding path or for solving a simple maze defined by aversive boundaries. This challenges the agent to build a world model that represents an extended spatial policy rather than a single boundary. See Listing 3.
    Define a maze as a grid where 1=path, 0=wall
    maze_layout = [
    [1, 1, 1, 0, 1],
    [0, 1, 0, 0, 1],
    [1, 1, 1, 1, 1],
    [1, 0, 1, 0, 1],
    [1, 1, 1, 0, 1]
    start_pos, goal_pos = (0, 0), (4, 4)
    agent_pos = start_pos
    while experiment_running:
    # State and Action decoding remain similar to 2D avoidance
    action = decode_4way_movement(groups_A_B)
    next_pos = calculate_next_pos(agent_pos, action)
    # 3. Feedback and state transition based on maze structure
    x, y = next_pos
    if maze_layout[y][x] == 1: # Valid path
    agent_pos = next_pos
    if agent_pos == goal_pos:
    deliver_reward(groups_C_D)
    agent_pos = start_pos # Reset to start
    else: # Hit a wall
    deliver_punishment(groups_C_D)
    # Agent does not move if it tries to enter a wall
    Listing 3: Pseudocode for maze navigation task
  • Dynamic Environments and Adaptation: To probe the adaptability of the learned world model, the environment can be made non-stationary. The boundary of the safe zone could slowly shift over the course of an experiment, or the goal location in a maze could change after a block of successful trials. This forces the agent to extinguish a previously learned policy and acquire a new one, providing a direct assay for the flexibility and update speed of its internal model (see Listing 4).
    Initialize environment with a moving boundary
    agent_position = 4
    safe_zone_boundary = 5 # Initially, positions 1-5 are safe
    boundary_shift_interval = 100 # Timesteps before boundary moves
    timestep = 0
    while experiment_running:
    # 1. State representation & 2. Action decoding (same as 1D)
    encode_position_as_stimulus(groups_C_D, agent_position)
    # 3. Feedback based on the DYNAMIC boundary
    if agent_position <= safe_zone_boundary:
    deliver_reward(groups_C_D)
    else:
    deliver_punishment(groups_C_D)
    # 4. Update the environment itself
    timestep += 1
    if timestep % boundary_shift_interval == 0:
    if random.choice([True, False]):
    safe_zone_boundary = max(1, safe_zone_boundary - 1)
    else:
    safe_zone_boundary = min(safe_zone_boundary + 1, 8)
    Listing 4: Pseudocode for dynamic avoidance task
4.2 Environment 2: Goal-seeking in a predator-prey scenario
4.2.1 Protocol description
  1. 1.
  2. 2.
  3. 3.
@@ 4.2.1 Protocol description
Listing 5: Pseudocode for predator-prey task
4.2.2 Environmental scaling and multi-agent systems
  • Dimensionality and Search Strategy: Scaling the world to 2D or 3D transforms the task from a simple directional problem into a genuine search problem. The agent must learn an efficient search strategy (e.g., patterned, memory-guided, or random) to locate the prey in a vastly larger state space. This places significant demands on the world model’s capacity to integrate sensory information over extended time horizons to infer the target’s location. See Listing 6.